# Mitigating Negative Side Effects in Multi-Agent Systems Using Blame Assignment

### v1

@@ Abstract
When agents that are independently trained (or designed) to complete their individual tasks are deployed in a shared environment, their joint actions may produce negative side effects (NSEs). As their training does not account for the behavior of other agents or their joint action effects on the environment, the agents have no prior knowledge of the NSEs of their actions. We model the problem of mitigating NSEs in a cooperative multi-agent system as a Lexicographic Decentralized Markov Decision Process with two objectives. The agents must optimize the completion of their assigned tasks while mitigating NSEs. We assume independence of transitions and rewards with respect to the agents’ tasks but the joint NSE penalty creates a form of dependence in this setting. To improve scalability, the joint NSE penalty is decomposed into individual penalties for each agent using credit assignment, which facilitates decentralized policy computation. Our results in simulation on three domains demonstrate the effectiveness and scalability of our approach in mitigating NSEs by updating the policies of a subset of agents in the system.
1 Introduction
Many real-world settings involve the operation of multiple autonomous agents that do not require coordination for the completion of their individual assigned tasks in an environment, such as robots with distinct areas of operation in factory warehouses [?]. Such agents are often trained in isolation and their training does not account for potential negative interactions with other agents or their cumulative effects on the environment. Consequently, the agents may act in ways that are collectively undesirable when deployed, such as producing negative side effects (NSEs) [??]. This paper focuses on mitigating NSEs in cooperative multi-agent settings where the agents produce no (or negligible) NSEs when executing their policy in isolation, but their joint policy results in NSEs.
@@ 1 Introduction
Consider warehouse robots, operating alongside humans, that optimize moving shelves between two locations. Each robot’s model provides the necessary information, including reward and transition dynamics, to complete its task optimally. The models may lack information about the effects of agents’ joint actions in the environment, such as a narrow corridor being blocked for human access when multiple robots simultaneously move large shelves through it. Thus, even when the agents are adept at completing their tasks and produce no NSEs when acting in isolation, their joint actions may have undesirable effects.
Figure 1: Overview of our solution approach. The agents independently compute optimal policies to complete their assigned tasks described by $R_{1}^{i}$ (Naive policy). The NSE Monitor computes the NSE penalty associated with the joint policy $\vec{\pi}$. The Blame Resolver then performs blame assignment to decompose the penalty in each joint state-action pair into penalties associated with each agent’s local state-action pair, which yields $R^{i}_{N}$. The agents finally recompute their policies by solving the bi-objective problem using lexicographic value iteration (LVI) with $R_{1}^{i}\succ R_{N}^{i}$.
@@
Mitigating NSEs in multi-agent settings is challenging because: (1) NSEs are often discovered after deployment, since they depend on the environment, agent interactions and their assigned tasks; (2) NSEs and corresponding penalties are defined over joint actions, introducing a dependency between agents; and (3) the computational complexity of mitigating NSEs, without significantly affecting task completion, increases with the number of agents in the system. Prior research on NSEs primarily target single-agent settings [????]. It is not straightforward to apply them to multi-agent settings as they do not account for the agent interactions that produce NSEs. Recently, [?] proposed an approach to avoid NSEs by optimizing the future returns by other agents in the environment. However, it cannot avoid NSEs that require policy update for multiple agents and do not fully scale to large settings.
@@
We formulate the problem of mitigating NSEs in a multi-agent system as a decentralized Markov decision process (DEC-MDP) with two objectives and a lexicographic ordering over them. The primary objective for each agent is to optimize its assigned task and the secondary objective is to minimize the NSEs. We assume independence of transition and rewards with respect to the agents’ assigned tasks [?]. Each agent’s reward for task completion is determined by its local state and actions and the overall reward for the system is the sum of individual agent’s rewards, and agents’ state transitions are affected only by their own actions. We use a lexicographic formulation since it is more intuitive and many real-world problems have an inherent ordering over objectives. It has been also been shown that lexicographic MDPs can be converted to a constrained MDP [?].
The agents do not have prior knowledge about NSEs but incur a joint penalty for them when executing certain actions together. Thus, while agents can independently compute policies to complete their tasks, mitigating NSEs requires addressing a form of dependence created by the joint penalty.
@@
We present a metareasoning [??] approach to detect and mitigate NSEs. The metareasoner is a centralized entity that monitors the agents’ behaviors and has two components: (1) NSE Monitor that estimates the NSEs associated with agents’ joint policy, and outputs the corresponding penalty, and (2) a Blame Resolver that decomposes the joint penalty into individual agent penalties, using our algorithm Reward Estimation using Counterfactual Neighbors (RECON). It is assumed that the NSE monitor has access to a model of NSEs and the associated penalty, either provided as part of its design or acquired through human feedback. The blame resolver, using RECON, performs blame (credit) assignment to determine each agent’s relative contribution towards the NSE, based on which the joint NSE penalty is decomposed into local penalties for each agent [???]. This decomposition facilitates updating the model of each agent with NSE information via a penalty function, thereby enabling decentralized policy computation to mitigate NSEs.
@@
Our solution framework uses a four-step approach to mitigate NSEs (Figure 1): (1) the agents first calculate optimal policies to complete their assigned tasks (referred to as naive policies); (2) the NSE monitor estimates the NSE penalty associated with the joint policy of the agents; (3) the blame resolver then decomposes the joint penalty into individual penalties for each agent, using blame (credit) assignment; and (4) the agents recompute their policies by solving a decentralized, bi-objective problem, with the prescribed reward for their task and the estimated local penalty for NSE, using lexicographic value iteration [?].
Our primary contributions are: (1) formalizing the problem of mitigating NSE in cooperative multi-agent settings as a decentralized, bi-objective problem; (2) presenting a metareasoning approach to update agent models with a penalty function corresponding to NSEs; (3) introducing an algorithm to decompose joint NSE penalty into individual penalties, using counterfactual-based credit assignment; and (4) empirical evaluation on three domains in simulation.
@@
Table 1: Overview of different credit assignment techniques for mitigating side effects. A technique is scalable if it is computationally inexpensive to generate counterfactuals for problems with a few hundreds of agents. “--” indicates that the approach cannot be applied directly but can be modified to meet the requirements of our setting.
2 Related Works
Negative Side Effects
Prior works have considered different definitions of NSEs such as hindering the completion of future tasks or the operation of other agents [????], and model incompleteness in the form of missing information or incomplete reward function [????]. Our setting is similar to the latter works in that the agents have incomplete information about the cumulative effects of their joint actions. A common approach in such settings is to modify the existing reward function or to consider a secondary reward function associated with NSEs [????] and solve a multi-objective problem. The information about NSEs may be gathered using human feedback [??] . However, these approaches have been presented for a single agent setting or consider other agents as part of the environment, and it is not straightforward to extend them to multi-agent settings where multiple agent policies must be updated. Recently, a distributed coordination approach has been proposed for mitigating NSEs in multi-agent settings [?]. They target NSEs arising due to local agent interactions and formulate it as a distributed constraint optimization problem and the agents learn a policy to avoid NSEs using coordinated Q-learning. Their approach is not scalable as distributed constraint optimization can become intractable for large problems [?]. We present a metareasoning approach that scales to large multi-agent systems.
Credit Assignment.
Credit assignment is a popular approach to measure the contribution of an agent to team performance, which is used to convert a joint reward into individual agent rewards [?]. Difference Reward [?] and its variants such as D++ [?], Wonderful Life Utility and Aristrocratic Utility [?], perform credit assignment by comparing the joint rewards before and after the agent is removed from the system. Another type of credit assignment uses Shapley value and involves computing a value function for each agent by considering all combinations of possible agent interactions [????]. Table 1 summarizes the characteristics of different credit assignment techniques. Blame assignment for NSEs must be based on counterfactuals generated by considering NSE-relevant features only. The existing methods calculate counterfactuals by considering all state features which leads to incorrect blame assignment for NSEs.
3 Preliminaries
Decentralized Markov Decision Process (Dec-MDP).
A Dec-MDP is widely used to model decentralized multi-agent decision-making problems [?]. A Dec-MDP is defined by the tuple $\langle\mathcal{A},S,A,T,R\rangle$ with $\mathcal{A}$ denoting the finite set of $k$ agents in the system; $S\!=\!\hat{S}_{1}\!\times...\times\!\hat{S}_{k}$ denoting the joint state space, where $\hat{S}_{i}$ denotes the state space of agent $i$; $A=\hat{A}_{1}\!\times\ldots\!\times\!\hat{A}_{k}$ denoting the joint action space, where $\hat{A}_{i}$ denotes agent $i$’s actions; $T:S\times A\times S\rightarrow[0,1]$ denoting the transition function; and $R$ denoting the reward function. A joint policy $\vec{\pi}\!=\!(\pi_{1},...,\pi_{k})$ is a set of policies, one for each agent in the system.
Transition and reward independence.   A Dec-MDP with transition independence and reward independence [?] is a class of problems in which agents operate independently but are tied together through a reward structure that depends on all of their execution histories. A transition and reward-independent MDP satisfies the following:
$T(\hat{s}_{i}^{\prime}|{s},{a},{s}^{\prime}_{-i})=T_{i}(\hat{s}_{i}^{\prime}|\hat{s}_{i},\hat{a}_{i}),\forall i\in\mathcal{A}\vspace{-3pt}$
@@ 3 Preliminaries
A transition and reward-independent Dec-MDP can be solved as $k$ single agent MDPs [?].
Lexicographic MDP (LMDP).
LMDPs are particularly convenient to model problems with potentially competing objectives with an inherent lexicographic ordering over them, such as ours where task completion is prioritized over NSE mitigation [??]. An LMDP is denoted by the tuple $M=\langle S,A,T,\boldsymbol{\mathrm{R}},o\rangle$ with finite set of states $S$, finite set of actions $A$, transition function denoted by $T\!:\!S\!\times\!A\times\!S\!\rightarrow\![0,1]$ and a vector of reward functions $\boldsymbol{\mathrm{R}}\!=\![R_{1},...,R_{k}]^{T}$ with $R_{i}\!:\!S\!\times\!A\!\rightarrow\!\mathbb{R}$, and $o$ denotes the strict preference ordering over the $k$ objectives. The set of value functions is denoted by $\boldsymbol{V}\!=\![V_{1},...,V_{k}]^{T}$, with $V_{i}$ corresponding to $o_{i}$,
@@
A slack $\boldsymbol{\Delta}\!=\!\langle\delta_{1},...,\delta_{k}\rangle$ with $\delta_{i}\!\geq\!0$, denotes the acceptable deviation from the optimal expected reward for objective $o_{i}$ so as to improve the lower priority objectives. Objectives are processed in the lexicographic order. The set of restricted actions for $o_{i+1}$ is $A_{i+1}(s)\!=\!\{a\!\in\!A|\max\limits_{a^{\prime}\in A_{i}}Q_{i}(s,a^{\prime})-Q_{i}(s,a)\leq\eta_{i}$ where $\eta_{i}\!=\!(1-\gamma)\delta_{i}$,$\gamma\in[0,1)$. Refer [?] for a detailed background on LMDP.
4 Problem Formulation
Problem Setting.  Consider a setting where $m$ agents are independently performing their assigned tasks which is their primary objective $o_{1}\!=\!\{o_{1}^{1},...,o_{1}^{m}\}$. The agents operate based on a transition and reward independent decentralized Markov decision process (Dec-MDP), denoted by $M$ [?]. A meta-level process, metareasoner, monitors and controls the agents performance (object-level process), by intervening if agent actions have undesirable consequences [??].
While the model $M$ contains all the information necessary to optimize $o_{1}$, it may not describe the joint effects of agents’ actions, since it is irrelevant to task completion in this setting. Due to the limited fidelity of $M$, negative side effects (NSEs) occur when agents execute their policies jointly. The agents incur a penalty for NSEs, determined by a function $R_{N}$ which is known to the metareasoner (either by design or learned using human feedback) but unknown to the agents.
We make the following assumptions about the NSEs: (1) there are no (or negligible) NSEs when agents execute their policy in isolation but their joint policy produces NSEs that must be mitigated to the extent possible; (2) agents have no prior knowledge about NSEs of the joint actions, besides receiving a penalty assigned by the metareasoner; and (3) NSEs do not interfere with the agents’ task completion—they are undesirable but are not catastrophic.We target settings where the completion of the agents’ assigned tasks ($o_{1}$) are prioritized over minimizing NSEs ($o_{2}$), $o_{1}\!\succ\!o_{2}$.
@@ 4 Problem Formulation
We consider independence of transition function and task completion reward $R_{1}$ in the MASE-MDP, meaning each agent’s transitions and task reward depend only on its local state and action. Hence, each agent can independently compute a policy for its assigned task by solving an MDPs [??]. However, the agents incur a joint penalty for the NSEs, which introduces a form of dependence and prevents decentralized computation of individual agent policies. In Section 5, we present an approach to decompose joint penalties into individual penalties, thereby facilitating decentralized planning.
Local and Global State Features.  We consider a factored state representation. Let $F$ denote the set of features in the environment, which are categorized into local features $F_{l}$ and global features $F_{g}$, $F\!=\!F_{l}\cup\!F_{g}$. The local features of an agent $i$ are denoted by $F^{i}_{l}$. Local features are agent-specific features that are controlled by the agent’s actions and affect its performance (e.g. an agent’s $x,y$ location). Global features are shared among agents and denote the overall state of the system. They are further divided into static and dynamic global features, denoted by $F_{gs}$ and $F_{gd}$ respectively, based on whether they are exogenous or can be modified by agent actions. An agent’s state $\hat{s}$ is described by $\vec{f}\!=\vec{f}_{l}^{i}\cup\vec{f}_{gd}\cup\vec{f}_{gs}$.
@@
Dynamic global features $F_{gd}$ describe properties of the environment that affect agent operation directly or indirectly, and can be modified by agent actions (e.g. locations and sizes of the shelves moved by the agents).
In this work, the penalty for NSE associated with a joint action is modeled as a function of the dynamic global features of the joint state, $R_{N}({s})=\Omega(\vec{f}_{gd})$. However, the proposed framework and approach can be extended to handle penalties defined over joint states and actions.
5 Penalty Decomposition via Blame Assignment

### v2

# Mitigating Side Effects in Multi-Agent Systems Using Blame Assignment

@@ Abstract
When independently trained or designed robots are deployed in a shared environment, their combined actions can lead to unintended negative side effects (NSEs). To ensure safe and efficient operation, robots must optimize task performance while minimizing the penalties associated with NSEs, balancing individual objectives with collective impact. We model the problem of mitigating NSEs in a cooperative multi-agent system as a bi-objective lexicographic decentralized Markov decision process. We assume independence of transitions and rewards with respect to the robots’ tasks, but the joint NSE penalty creates a form of dependence in this setting. To improve scalability, the joint NSE penalty is decomposed into individual penalties for each robot using credit assignment, which facilitates decentralized policy computation. We empirically demonstrate, using mobile robots and in simulation, the effectiveness and scalability of our approach in mitigating NSEs.
I INTRODUCTION
In many real-world environments, the actions of individual robots, even when operating independently, can have significant and often unintended consequences on the broader environment and the collective system behavior. Traditional approaches to agent design and training focus on optimizing performance in isolation, prioritizing task completion while overlooking potentially harmful effects their actions can collectively produce, such as negative side effects (NSEs). In multi-agent systems, NSEs are the unintended and undesirable outcomes of collective system behavior, caused by the incompleteness of models used for decision making [1, 20, 21]. While the NSEs are difficult to identify before deployment, mitigating them is crucial for overall system efficiency and safety.
@@ I INTRODUCTION
This paper focuses on mitigating NSEs in cooperative multi-agent settings where the robots produce no (or negligible) NSEs when executing their policy in isolation, but their joint policy results in NSEs. Consider warehouse robots that optimize moving shelves between two locations. Each robot’s model provides the necessary information, including reward and transition dynamics, to complete its task optimally. The models lack information about the effects of robots’ joint actions in the environment, such as a narrow corridor being blocked for human access when multiple robots simultaneously move large shelves through it. Thus, even when the robots are adept at their tasks and produce no NSEs when acting in isolation, their joint actions are undesirable.
Figure 1: Comparison of the paths taken by two TurtleBots performing delivery tasks in our indoor setup. The robots, initially unaware about the side effects of their actions, receive a joint penalty when one or both are in NSE states marked by X. The robots must update their behavior to complete tasks while mitigating NSEs.
@@
Mitigating NSEs in multi-agent settings is challenging because NSEs and corresponding penalties are defined over joint actions, creating inter-agent dependencies that did not exist for task completion. Further, the computational complexity of mitigating NSEs, without significantly affecting task completion, increases with the number of agents in the system. Prior research on NSEs primarily target single-agent settings [10, 11, 20, 21, 29] or treat other agents as part of the environment [1]. These techniques do not apply to multi-agent settings as they ignore the agent interactions that produce NSEs. A recent approach uses distributed constraint optimization with Q-learning to mitigate multi-agent NSEs [6]. However, it is not scalable since distributed constraint optimization is intractable for large problems [7].
Figure 2: Overview of our solution approach. Agents independently compute policies to complete tasks described by $R_{1}^{i}$ (Naive policy). The NSE Monitor computes the NSE penalty for the joint policy $\vec{\pi}$. The Blame Resolver assigns a blame value for each agent, by evaluating counterfactual scenarios specific to each agent, as illustrated with warehouse robots handling different-sized boxes. Individual penalty functions $R^{i}_{N}$ are derived for each agent, based on the estimate blame. Agents then recompute their policies by solving the bi-objective problem with $R_{1}^{i}\succ R_{N}^{i}$, where $\succ$ denotes preference ordering over the objectives and their associated reward functions.
@@
We formulate the problem of mitigating NSEs in a multi-agent system as a decentralized Markov decision process (DEC-MDP) with two objectives and a lexicographic ordering over them. A lexicographic formulation is an intuitive way to model problems with an inherent ordering over objectives [19, 27]. The primary objective for each agent is to optimize its assigned task. We assume independence of transition and rewards with respect to the agents’ assigned tasks [2]. That is, each agent’s reward for task completion is determined by its local state and actions and the overall reward for the system is the sum of individual agent’s rewards, and agents’ state transitions are affected only by their own actions. The secondary objective is to minimize the NSEs. Agents have no prior knowledge of NSEs and only observe the joint penalty—a global value based on all agents’ actions. While they can independently compute policies for their tasks, mitigating NSEs requires addressing the dependence induced by the joint penalty (Fig. 1).
@@
We present a metareasoning [4, 32] approach to detect and mitigate NSEs. The metareasoner is a centralized entity that monitors the agents’ behaviors and has two components: (1) NSE Monitor that estimates the NSEs associated with agents’ joint policy, and outputs the corresponding penalty, and (2) a Blame Resolver that decomposes the joint penalty into individual agent penalties, using our algorithm Reward Estimation using Counterfactual Neighbors (RECON). It is assumed that the NSE monitor has access to a model of NSEs and the associated penalty, either provided as part of its design or acquired through human feedback. The blame resolver, using RECON, performs blame (credit) assignment to determine each agent’s relative contribution towards the NSE, based on which the joint NSE penalty is decomposed into local penalties for each agent [3, 15, 17]. This decomposition facilitates augmenting the model of each agent with NSE information via a penalty function, thereby enabling decentralized policy computation to mitigate NSEs.
@@
Our solution framework uses a four-step approach to mitigate NSEs (Fig. 2): (1) the agents first calculate optimal policies to complete their assigned tasks (referred to as naive policies); (2) the NSE monitor estimates the NSE penalty associated with the joint policy of the agents; (3) the blame resolver then decomposes the joint penalty into individual penalties for each agent, using blame (credit) assignment; and (4) the agents recompute their policies by solving a decentralized, bi-objective problem, with the prescribed reward for their task and the estimated local penalty for NSE, using lexicographic value iteration [27]. Our experiments using mobile robots and in simulation demonstrate the efficiency and scalability of our approach in mitigating NSEs by updating the policies of a subset of agents in the system.
TABLE I: Overview of different credit assignment techniques for mitigating NSEs. A technique is scalable if it is computationally inexpensive to generate counterfactuals for problems with a few hundreds of agents. “--” indicates that the approach cannot be applied directly but can be modified to meet the requirements of our setting.
@@
II BACKGROUND
Decentralized Markov Decision Processes (Dec-MDPs)  are widely used to model decentralized multi-agent decision-making problems [9]. A Dec-MDP is defined by the tuple $\langle\mathcal{A},S,A,T,R\rangle$ with $\mathcal{A}$ denoting the finite set of $k$ agents in the system; $S\!=\!\hat{S}_{1}\!\times\!\ldots\!\times\!\hat{S}_{k}$ denoting the joint state space, where $\hat{S}_{i}$ denotes the state space of agent $i$; $A=\hat{A}_{1}\!\times\ldots\!\times\!\hat{A}_{k}$ denoting the joint action space, where $\hat{A}_{i}$ denotes agent $i$’s actions; $T:S\times A\times S\rightarrow[0,1]$ denoting the transition function; and $R$ denoting the reward function. We use $s_{-i}$ to denote joint state excluding agent $i$. A joint policy $\vec{\pi}\!=\!(\pi_{1},...,\pi_{k})$ is a set of policies, one for each agent in the system. A Dec-MDP with transition independence and reward independence [2] is a class of problems in which agents operate independently but are tied together through a reward structure that depends on all of their execution histories:
$T(\hat{s}_{i}^{\prime}|{s},{a},{s}^{\prime}_{-i})=T_{i}(\hat{s}_{i}^{\prime}|\hat{s}_{i},\hat{a}_{i}),\forall i\in\mathcal{A},$
@@ II BACKGROUND
A transition and reward-independent Dec-MDP can be solved as $k$ single agent MDPs [9].
Lexicographic MDP (LMDP)  LMDPs are particularly convenient to model problems with potentially competing objectives and an inherent lexicographic ordering over them, such as ours where task completion is prioritized over NSE mitigation [20, 27]. An LMDP is denoted by the tuple $M=\langle S,A,T,\boldsymbol{\mathrm{R}},o\rangle$ with finite set of states $S$, finite set of actions $A$, transition function denoted by $T\!:\!S\!\times\!A\times\!S\!\rightarrow\![0,1]$ and a vector of reward functions $\boldsymbol{\mathrm{R}}\!=\![R_{1},...,R_{k}]^{T}$ with $R_{i}\!:\!S\!\times\!A\!\rightarrow\!\mathbb{R}$, and $o$ denotes the strict preference ordering over the $k$ objectives. The set of value functions is denoted by $\boldsymbol{V}\!=\![V_{1},...,V_{k}]^{T}$, with $V_{i}$ corresponding to $o_{i}$,
@@
A slack $\boldsymbol{\Delta}\!=\!\langle\delta_{1},...,\delta_{k}\rangle$ with $\delta_{i}\!\geq\!0$, denotes the acceptable deviation from the optimal expected reward for objective $o_{i}$ so as to improve the lower priority objectives. Objectives are processed in the lexicographic order. The set of restricted actions for $o_{i+1}$ is $A_{i+1}(s)\!=\!\{a\!\in\!A|\max\limits_{a^{\prime}\in A_{i}}Q_{i}(s,a^{\prime})-Q_{i}(s,a)\leq\eta_{i}\}$, where $\eta_{i}\!=\!(1-\gamma)\delta_{i}$,$\gamma\in[0,1)$. Refer [27] for a detailed background on LMDP.
Credit Assignment  It is a popular approach to measure the contribution of an agent to team performance so as to convert a joint reward into individual agent rewards [15]. Difference Reward [17] and its variants such as D++ [18], Wonderful Life Utility [15], perform credit assignment by comparing the joint rewards before and after the agent is removed from the system. Another type of credit assignment uses Shapley value to compute a value function for each agent by considering all combinations of possible agent interactions [3, 14, 22, 23]. Table I summarizes the characteristics of different credit assignment techniques. While some of these techniques assess individual contributions, they do not support decentralized planning due to their reliance on other agents’ policies. The existing methods calculate counterfactuals by considering all state features, including those that are not associated with NSEs, which leads to incorrect blame assignment for NSEs.
III PROBLEM FORMULATION
Problem Setting.  Consider $m$ agents independently performing their assigned tasks which is their primary objective $o_{1}\!=\!\{o_{1}^{1},...,o_{1}^{m}\}$. Each agent operates based on an MDP $\hat{M}$ that contains all the information necessary to optimize $o_{1}$ but lacks information about the joint effects of agents’ actions when it is unrelated to task completion. Due to the limited fidelity of $\hat{M}$, negative side effects (NSEs) occur when agents execute their policies simultaneously. A meta-level process, metareasoner, monitors and controls the agents’ performance (object-level process) [26, 32], by providing a penalty if agents’ actions produce side effects. The NSE penalty is determined by a function $R_{N}$ which is known to the metareasoner (either by design or learned using human feedback) but unknown to the agents.
We assume the following about the NSEs: (1) agents acting in isolation produce no (or negligible) NSEs, but their joint actions produce NSEs that must be mitigated; (2) agents have no prior knowledge about NSEs of the joint actions except for the penalty assigned by the metareasoner; and (3) NSEs are undesirable but do not hinder task completion. We address the problem of mitigating NSEs by augmenting the agents’ models with secondary reward functions that correspond to NSE penalties. We target settings where the completion of the agents’ assigned tasks ($o_{1}$) are prioritized over minimizing NSEs ($o_{2}$), $o_{1}\!\succ\!o_{2}$.
@@ III PROBLEM FORMULATION
A MASE-MDP is characterized by independence of transition function and task completion reward $R_{1}$, meaning each agent’s transitions and task reward depend only on its local state and action and is independent of other agents. This allows each agent to independently compute a policy for its assigned task [2, 9]. However, the agents incur a joint penalty for the NSEs, which introduces a form of inter-agent dependence and prevents decentralized computation of individual policies. Section IV describes an approach to decompose joint penalties into individual penalties, thereby facilitating decentralized planning.
Local and Global State Features.  We consider a factored state representation. Let $F$ denote the set of features in the environment, which are categorized into local features $F_{l}$ and global features $F_{g}$, $F\!=\!F_{l}\cup\!F_{g}$. The local features of an agent $i$ are denoted by $F^{i}_{l}$. Local features are agent-specific features that are controlled by the agent’s actions and affect its performance (e.g. $x,y$ location). Global features are shared among agents and denote the overall state of the system. They are further divided into static and dynamic global features, denoted by $F_{gs}$ and $F_{gd}$ respectively, based on whether they are immutable or can be modified by agent actions. An agent’s state $\hat{s}$ is described by $\vec{f}\!=\vec{f}_{l}^{i}\cup\vec{f}_{gd}\cup\vec{f}_{gs}$.
@@
Dynamic global features $F_{gd}$ describe properties of the environment that affect agent operation directly or indirectly, and can be modified by agent’s actions (e.g. locations and sizes of the shelves moved by the agents).
We model the NSE penalty as a function of dynamic global features of the joint state, $R_{N}({s})\!=\!\Omega(\vec{f}_{gd})$, where $\Omega$ is a mapping from $\vec{f}_{gd}$ to the penalty value. This can also be written in the common state-action notation as $R_{N}(s,a)\!=\!R_{N}(s),\forall a\!\in\!A$.
IV PENALTY DECOMPOSITION